OpenStudy (anonymous):

Two regular pentagons have perimeters of 8cm and 16cm. What is the ratio of their areas?

OpenStudy (anonymous):

1:2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

perimeter is a linear measure, the scale factor would be 1:2 however, you'd have to square these numbers to get the scale factor for the areas

OpenStudy (anonymous):

i don't quite understand sorry

OpenStudy (anonymous):

let's look at something simple, like a square.  do you see one of the inside squares? the outside square has twice the length. but how many time the area?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

yep, \(4=2^2\)

OpenStudy (anonymous):

is it 8^2 / 16^2 ? which equals to 4

OpenStudy (anonymous):

same for circles, too. if a smaller circle has a radius of 3, it will have an area of \(\pi r^2 = 9\pi\) if you double the radius to 6, the area is now \(\pi r^2 = 36\pi\) which is 4 times more.

OpenStudy (anonymous):

Oh ok I see now. I didn't know it asked for area

OpenStudy (anonymous):

that works. if \(P_1:P_2 = 1:2 \Rightarrow P_1^2:P_2^2 = 1^2:2^2 = 1:4\) the P is the perimeter, a linear measure. for volumes, you have to cube the linear scale factor.

OpenStudy (anonymous):

you good?

OpenStudy (anonymous):

i do understand a bit but how does that get me the ratio of areas ?

OpenStudy (anonymous):

how did you know the perimeter of p1 and p2 is 1:2

OpenStudy (anonymous):

it's given... 8:16 = 1:2

OpenStudy (anonymous):

OHH ok because they both are multiples of 8 sorry im quite slow at this haha

OpenStudy (anonymous):

also, \(8:16=\Large{\frac{8}{16}}\normalsize{=}\Large{\frac{1}{2}}\)

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

is 1:2 the finally anwser ?

OpenStudy (anonymous):

no... you have to square those numbers to getthe ratio of the areas. the 1:2 is for their perimeters (a linear measure)

OpenStudy (anonymous):

If it was a line you would be using 1:2 but because it is area that they are asking for you have to square it \[1^2:2^2\]

OpenStudy (anonymous):

@pgpilot326 am I right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so 1:4 ?

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

thankds !

OpenStudy (anonymous):

you're welcome!